Angle U and angle V are complemetary angles. If Angle U is three less than twice angle V, What are the measurements of the angles?
Answer by IWork4Dessert(60)

Complementary angles are angles whose values, when added together, equal 90º.
Let angle V equal "x" and angle U equal "2x-3", as the instructions state that angle U is three less than twice angle V.
Now you create an equation using this.
(x)+(2x-3)=90
3x-3=90
+3 +3
3x=93
x=31
Plug in the answer to your variables.
x=31
2x-3=59
To check your answer, add them together and make sure they equal 90º.
31+59=90
So your final answer is:
Angle U=59º
Angle V=31º
Hope this helps!